OpenStudy (anonymous):

How do I "find all primitive pythagorean triples of the form (40, y, z)" ? I know (40, 9, 41) is a pythagorean triple, but I'm sure how to find the solution for all of them.

OpenStudy (anonymous):

Thanks, but I think it's a bit different from that question. I think the answer might be a general form.

OpenStudy (anonymous):

right ... didn't read all primitives

OpenStudy (anonymous):

Ok :)

OpenStudy (anonymous):

a= k (m^2-n^2), b=k (2mn), c = k ( m^2 +n^2 )

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Pythagorean_triple

OpenStudy (anonymous):

Oh, I see! Thank you! :)